# Optimal control of stochastic reaction networks with entropic control cost and emergence of mode-switching strategies

### v1

@@ Abstract
Controlling the stochastic dynamics of biological populations is a challenge that arises across various biological contexts. However, these dynamics are inherently nonlinear and involve a discrete state space, i.e., the number of molecules, cells, or organisms. Additionally, the possibility of extinction has a significant impact on both the dynamics and control strategies, particularly when the population size is small. These factors hamper the direct application of conventional control theories to biological systems. To address these challenges, we formulate the optimal control problem for stochastic population dynamics by utilizing a control cost function based on the Kullback–Leibler divergence. This approach naturally accounts for population-specific factors and simplifies the complex nonlinear Hamilton–Jacobi–Bellman equation into a linear form, facilitating efficient computation of optimal solutions. We demonstrate the effectiveness of our approach by applying it to the control of interacting random walkers, Moran processes, and SIR models, and observe the mode-switching phenomena in the control strategies. Our approach provides new opportunities for applying control theory to a wide range of biological problems.
@@ I Introduction
Optimal control problems for a population of stochastically interacting particles arise in diverse fields of biology [1, 2]. In intracellular chemical reactions, molecules interact stochastically and nonlinearly to generate complex dynamics, whose control is essential for medical and bioengineering applications [3]. In cellular or animal populations, cells or organisms with different phenotypic and genetic traits interact and compete for survival and reproduction. Strategic control of such populations into either extinction, survival, or amplification leads to cancer therapy [4, 5, 6], stem cell culturing [7], manipulating gut microbiota [8], maintenance of immunological memory [9, 10], biodiversity conservation [11, 12, 13], and control of evolving population [14, 15] In human populations, control of epidemic outbreaks, spurred by individual interactions, poses a significant public health challenge [16, 17, 18].
@@
Despite its broad applicability, the optimal control of RNs remains underexplored. This oversight is largely due to the unique characteristics of stochastic RNs, which deviate from the setup of conventional optimal control for diffusion processes. The major deviations include the inherent nonlinearity of RNs, the discrete nature of state variables (representing counts of particles like molecules or organisms), and their stochastic dynamics, which are better modeled by Markov jump processes with Poissonian randomness rather than Gaussian diffusion [19, 1, 20]. Additionally, the non-negativity constraint of control parameters, i.e., kinetic rate constants, requires a natural measure for cost other than conventional quadratic ones, which presume the constraint-free Euclidian parameter space. Finally, the zero count states act as intrinsic absorbing boundaries, reaching them can dramatically alter system dynamics through the extinction of particles (Fig. 1).
These distinct properties of RNs necessitate (1) a departure from the traditional control theories based on diffusion processes, such as the Linear-Quadratic-Gaussian (LQG) model, (2) an establishment of alternative theoretical framework tailored to the unique requirements of RN optimal control, and (3) its applications to biologically relevant control problems.
In this work, we establish that all these issues can be resolved by integrating optimal control of jump processes and stochastic RN with a cost function based on relative entropy or Kullback–Leibler (KL) divergence. The optimal control with KL cost was originally proposed for diffusion processes in relation to the duality of control and inference [21, 22, 23]. We demonstrate that KL cost, as it is designed for non-negative probabilities and densities, can naturally accommodate the non-negativity constraints for both state space (particle counts) and control parameters (kinetic rate constants) of stochastic RN. Moreover, KL cost allows for the linearization of the nonlinear Hamilton–Jacobi–Bellman (HJB) equations through the Cole-Hopf transformation. This linearization facilitates the efficient derivation of optimal solutions in a similar matter with previously identified solvable control problems [24, 25, 26].
@@
All these results are obtained without directly solving nonlinear HJB equations, which are typically intractable, even numerically. Thus, our new results can substantially broaden the scope of optimal control applications in stochastic chemical reactions, population dynamics, and epidemics.
Figure 1: Examples of biological phenomena described as reaction networks. (a) Movement of molecular motors on a microtubule with mutual interference. (b) Competing dynamics of populations in population genetics or ecology. (c) Spread of infectious diseases in a population. Purple thick lines represent the absorbing states in each phenomenon.
@@ II.1 Stochastic reaction systems
The timing of reaction events is random, which follow an inhomogeneous Poisson process with rate $\lambda_{r}(t)\geq 0$ for reaction $r$ at time $t$. It could vary over time and depends on the current number distribution $n(t)$. We write the rate of reaction $r\in R$ as the product of the reaction rate coefficient $k_{r}(t)\geq 0$ and the propensity function $h_{r}(n)\geq 0$:
$\lambda_{r}(t)=k_{r}(t)h_{r}(n(t)).$ (2)
When we assume the law of mass action kinetics, $k_{r}(t)$ is a kinetic constant, and the function $h_{r}(n)$ is given by $h_{r}(n)=N^{1-\sum_{x\in X}s^{-}_{r,x}}\prod_{x\in X}\frac{n_{x}!}{(n_{x}-s^{-}{r,x})!}$, where $N$ is a constant parameter that describe the system size, e.g., the volume or total number of the particles in the system. We do not assume the mass action kinetics in the following so that our results become applicable to any propensity function. Instead, we presume only the situations that make the process well-defined up to time $T$ or up to exit time $T_{\mathrm{exit}}$. For example, $h_{r}(n)$ should be 00 when the reaction pushes the state out of the domain, i.e., $n+s_{r}\not\in\mathcal{N}$.
The stochastic reaction system defined in this way can be characterized by either the counting process representation or the Markov process representation. The counting process representation [20] is useful for characterizing the stochastic process $(n(t))_{0\leq t\leq T}$ and simplifies the derivation of KL control cost. Let $\xi_{r}(t)\in\mathbb{Z}_{\geq 0}$ denote the number of occurrences of reaction $r$ from time 00 to $t$. The number distribution $n(t)$ is a linear transformation of the reaction count $\xi(t)$, i.e., for any $0\leq t\leq T$,
@@
[…]
The stochastic process $n(\cdot)$ is a Markov process on the nonnegative integer lattice $\mathcal{N}$. The transition rate from $n\in\mathcal{N}$ to $n^{\prime}\in\mathcal{N}$ is given by
@@
$\frac{\partial}{\partial t}P_{t}(n)$ $=\sum_{n^{\prime}\in\mathcal{N}\backslash\{n\}}\left[P_{t}(n^{\prime})\omega_{n^{\prime},n}(t)-P_{t}(n)\omega_{n,n^{\prime}}(t)\right]$
@@
$=:\mathcal{L}_{k}^{\dagger}P_{t}(n)$ (5)
II.2 General formulation of optimal control problems
[…]
$\mathcal{E}(n_{0},u):={}$ $\underset{k(\cdot)}{\text{min}}~{}~{}\mathbb{E}_{k(\cdot)}\left[\mathcal{C}(n(\cdot),k(\cdot))\right]$ (6)
[…]
or equivalently, the following unconstrained optimization:
@@ II.2 General formulation of optimal control problems
where the expectation $\mathbb{E}_{k(\cdot)}$ is taken over trajectories $n(\cdot)$ generated under the designated reaction rate coefficients $k(\cdot)$ and initial condition $n(0)=n_{0}$. A scalar $\beta$ is the Lagrange multiplier or a parameter to adjust the importance of the utility relative to the control cost.
We will calculate the optimal transition rate $(k^{\dagger}(t))_{0\leq t<T}$ that attains the optimum of $\mathcal{E}(n_{0},u)$ and $V(n_{0},\beta)$. The standard protocol for solving the problem is to consider the value function $V_{t}(n,\beta)$, which is defined as the maximum of the expectation in Eq. (7) from $t$ to $T$ under the condition $n(t)=n$:
@@
[…]
If $V(n,\beta)$ were analytically or numerically obtained for all $n$ and $\beta$, $\mathcal{E}(n,u)$ in Eq. (6) is derived as the Legendre transformation of $V(n,\beta)$: (see Sec. A in the Supplementary Material  [^1] See Supplemental Material at for the derivation of equations and detailed numerical results, which includes Refs. [59, 60, 61, 62, 63, 64]. for the derivation)
$\mathcal{E}(n,u)=\max_{\beta}\left[u\beta-V(n,\beta)\right].$ (9)
Furthermore, the expected utility under the optimal control can also be derived from the value function (see Sec. A of the Supplementary Material [27] for the derivation)
$\mathcal{U}^{\dagger}(n,\beta):=\mathbb{E}_{k^{\dagger}(\cdot)}\left[\mathcal{U}(n(\cdot))\right]=\frac{\partial}{\partial\beta}V(n,\beta).$ (10)
II.3 Optimal control with KL cost
@@
To obtain the value function, we usually leverage the Hamilton–Jacobi–Bellman (HJB) equation, a differential equation that $V_{t}(n,\beta)$ satisfies. However, solving the HJB equation is generally intractable both analytically and even numerically because it is a nonlinear differential equation on a possibly infinite domain. This difficulty is the major obstacle to optimal control in applications. To address the difficulty, we are conventionally forced to restrict or approximate the original problem to a linear quadratic Gaussian (LQG) problem, in which only the linear dynamics on continuous Euclidean space with Gaussian noise with a quadratic control cost function can be considered. However, all these restrictions of the LQG problem conflict with the essential properties of stochastic RN control, i.e., the nonlinearity of $h_{r}(n)$, the discreteness and nonnegativity of state $n\in\mathbb{Z}_{\geq 0}^{|X|}$, the nonnegativity of the control parameter $k_{r}\geq 0$, and Poissonian nature of stochasticity. Several studies attempted to overcome a part of these difficulties [28, 29, 30]. Nonetheless, optimal control was not practical for biological problems described by RN.
In this work, we clarify that the difficulty can be resolved by adopting the following cost function
$C(n,k)=\sum_{r\in R}h_{r}(n)c_{KL}(k_{r},k^{0}_{r}),$ (11)
where $c_{KL}(a,a^{0})$ is the generalized Kullback–Leibler (KL) divergence for positive scalars $a,a^{0}>0$:
$c_{KL}(a,a^{0}):=a\log\frac{a}{a^{0}}-a+a^{0}.$ (12)
The value at $a=0$ is defined by the limit $c_{KL}(0,a^{0}):=\lim_{a\rightarrow 0}c_{KL}(a,a^{0})=a^{0}$.
The KL control cost function in Eq. (11) measures the deviation of the controlled reaction rate coefficient $k$ from its uncontrolled reaction rate coefficient $k^{0}\in\mathbb{R}^{|R|}_{>0}$. Because the control cost function $C(n,k)$ is minimized at $k=k^{0}$, the optimal reaction rate coefficient $k^{\dagger}$ is equal to $k^{0}$ when $\beta=0$, i.e., no preference on the state trajectory. If $\beta\neq 0$, the optimal reaction coefficient for $r\in R$ at time $t$ is given by

### v2

@@ Abstract
Controlling the stochastic dynamics of biological populations is a challenge that arises across various biological contexts. However, these dynamics are inherently nonlinear and involve a discrete state space, i.e., the number of molecules, cells, or organisms. Additionally, the possibility of extinction has a significant impact on both dynamics and control strategies, particularly when the population size is small. These factors hamper the direct application of conventional control theories to biological systems. To address these challenges, we formulate the optimal control problem for stochastic population dynamics by utilizing control cost functions based on the $f$-divergence, which naturally accounts for population-specific factors. If Kullback–Leibler (KL) divergence is adopted for the cost function, the complex nonlinear Hamilton–Jacobi–Bellman equation is simplified into a linear form, facilitating efficient computation of optimal solutions. We demonstrate the effectiveness of our approach by applying it to the control of interacting random walkers, Moran processes, and SIR models, and observe the mode-switching phenomena in the control strategies. Our approach provides new opportunities for applying control theory to a wide range of biological problems.
@@ I Introduction
Optimal control problems for a population of stochastically interacting particles arise in diverse fields of biology [1, 2]. In intracellular chemical reactions, molecules interact stochastically and nonlinearly to generate complex dynamics, whose control is essential for medical and bioengineering applications [3]. In cellular or animal populations, cells or organisms with different phenotypic and genetic traits interact and compete for survival and reproduction. Strategic control of such populations into either extinction, survival, or amplification leads to cancer therapy [4, 5, 6], stem cell culturing [7], manipulating gut microbiota [8], maintenance of immunological memory [9, 10], biodiversity conservation [11, 12, 13], and control of evolving population [14, 15] In human populations, control of epidemic outbreaks, spurred by individual interactions, poses a significant public health challenge [16, 17, 18, 19].
@@
Despite its broad applicability, the optimal control of RNs remains underexplored. This oversight is largely due to the unique characteristics of stochastic RNs, which deviate from the setup of conventional optimal control for diffusion processes. The major deviations include the inherent nonlinearity of RNs, the discrete nature of state variables (representing counts of particles like molecules or organisms), and their stochastic dynamics, which are better modeled by Markov jump processes with Poissonian randomness rather than Gaussian diffusion [20, 1, 21]. Additionally, the non-negativity constraint of control parameters, i.e., kinetic rate constants, requires a natural measure for cost other than conventional quadratic ones, which presume the constraint-free Euclidian parameter space. Finally, the zero count states act as intrinsic absorbing boundaries, reaching them can dramatically alter system’s dynamics through the extinction of particles (Fig. 1).
These distinct properties of RNs necessitate (1) a departure from the common control setup based on diffusion processes, such as the Linear-Quadratic-Gaussian (LQG) model, (2) an establishment of alternative theoretical framework tailored to the unique requirements of RN optimal control, and (3) its applications to biologically relevant control problems.
In this work, we establish that all these issues can be resolved by integrating optimal control of jump processes and stochastic RN with cost functions based on $f$-divergence [22, 23, 24], which includes relative entropy or Kullback–Leibler (KL) divergence. The optimal control with KL cost was originally proposed for diffusion processes in relation to the duality of control and inference [25, 26, 27]. We demonstrate that $f$-divergence, as it is designed for non-negative probability measures, can naturally accommodate the non-negativity constraints and the nature of stochastic RNs as counting processes. Moreover, KL cost allows for the linearization of the nonlinear Hamilton–Jacobi–Bellman (HJB) equations through the Cole-Hopf transformation. This linearization facilitates the efficient derivation of optimal solutions in a similar manner to previously identified solvable control problems [28, 29, 30].
@@
Thus, our new results can substantially broaden the scope of optimal control applications in stochastic chemical reactions, population dynamics, and epidemics.
Figure 1: Examples of biological phenomena described as reaction networks. (a) Movement of molecular motors on a microtubule with mutual interference. (b) Competing dynamics of populations in population genetics or ecology. (c) Spread of infectious diseases in a population. Purple thick lines in the panels represent the absorbing states in each phenomenon.
@@ II.1 Stochastic reaction systems
The timing of reaction events is random, which follow an inhomogeneous Poisson process with rate $j_{r}(t)\geq 0$ for reaction $r$ at time $t$. It could vary over time and depends on the current number distribution $n(t)$. We write the rate of reaction $r\in R$, or the reaction flux, as the product of the reaction rate coefficient $k_{r}(t)\geq 0$ and the propensity function $h_{r}(n)\geq 0$:
$j_{r}(t)=k_{r}(t)h_{r}(n(t)).$ (2)
When we assume the law of mass action kinetics, $k_{r}(t)$ is a kinetic rate constant, and the function $h_{r}(n)$ is given by $h_{r}(n)=N^{1-\sum_{x\in X}s^{-}_{r,x}}\prod_{x\in X}\frac{n_{x}!}{(n_{x}-s^{-}_{r,x})!}$, where $N$ is a constant parameter that describe the system size, e.g., the volume or total number of the particles in the system. We do not assume the mass action kinetics in the following so that our results become applicable to any propensity function. Instead, we presume only the situations that make the process well-defined up to time $T$ or up to exit time $T_{\mathrm{exit}}$. For example, $h_{r}(n)$ should be 0 when the reaction pushes the state out of the domain, i.e., $n+s_{r}\not\in\mathcal{N}$.
The stochastic reaction system defined in this way can be characterized by either the counting process representation or the Markov process representation. The counting process representation [21] is useful for characterizing the stochastic process $(n(t))_{0\leq t\leq T}$ and simplifies the derivation of KL control cost. Let $\xi_{r}(t)\in\mathbb{Z}_{\geq 0}$ denote the total number of occurrences of reaction $r$ from time 0 to $t$. The number distribution $n(t)$ is a linear transformation of the reaction count $\xi(t)$, i.e., for any $0\leq t\leq T$,
@@
The probability law of $\xi(\cdot)$ is given by the Poisson process. Using the time-change of independent unit Poisson processes $Y_{r}(t)$, the reaction count can be written as $\xi_{r}(t)=Y_{r}\left(\int_{0}^{t}k_{r}(\tau)h_{r}(n(\tau))d\tau\right)$.
The stochastic process $n(\cdot)$ is a Markov process defined on the nonnegative and infinite integer lattice $\mathcal{N}$. The transition rate from $n\in\mathcal{N}$ to $n^{\prime}\in\mathcal{N}$ is given by
@@
[…]
@@
$=:\mathcal{L}_{k}^{\dagger}P_{t}(n).$ (5)
II.2 General formulation of optimal control problems for stochastic RNs
We assume that the controller can observe the current state $n(t)$ and adjust the reaction rate coefficients $k_{r}(t)\geq 0$ to any desired value at any time while the function $h_{r}(n)$ remains unchanged, i.e., the controller can modulate only the speeds of reactions. This is a feedback control problem so that the reaction rate coefficients $k_{r}(t)$ can depend on the current state $n(t)$. We would like to find the optimal reaction rate coefficients $(k(t))_{0\leq t<T}$ which drive the population to take a desirable trajectory $(n(t))_{0\leq t<T}$ while minimizing the cost of modulating the reaction rate coefficients. Let us define the utility function $\mathcal{U}(n(\cdot)):=U_{T}(n(T))+\int_{0}^{T}U_{\tau}(n(\tau))d\tau$ as the sum of the terminal utility function $U_{T}(n)\in\mathbb{R}$ and the time integral of instantaneous utility function $U_{\tau}(n)\in\mathbb{R}$. The control cost function is also defined as $\mathcal{C}(n(\cdot),k(\cdot)):=\int_{0}^{T}C(n(\tau),k(\tau))d\tau$. Then, we consider the following constrained optimization problem:
$\mathcal{E}(n_{0},u)$= $\underset{k(\cdot)}{\text{min}}~~\mathbb{E}_{k(\cdot)}\left[\mathcal{C}(n(\cdot),k(\cdot))\right]$ (6)
$\text{s.\,t.}~~\mathbb{E}_{k(\cdot)}\left[\mathcal{U}(n(\cdot))\right]\geq u,$
or corresponding Lagrange dual problem [31], which is given by the following unconstrained optimization:
@@ II.2 General formulation of optimal control problems for stochastic RNs
where the expectation $\mathbb{E}_{k(\cdot)}$ is taken over trajectories $n(\cdot)$ generated under the designated reaction rate coefficients $k(\cdot)$ and initial condition $n(0)=n_{0}$. A scalar $\beta$ is the Lagrange multiplier or a parameter to adjust the importance of the utility relative to the control cost. Once we could find the optimal reaction rate coefficient $(k^{\dagger}(t))_{0\leq t<T}$ that solves one of these problems for all $u$ or for all $\beta$ such that the solution exists, the solution to the other problem is readily obtained by the Legendre transformation:
$\mathcal{E}(n,u)=\max_{\beta}\left[u\beta-V(n,\beta)\right],$ (8)
$V(n,\beta)=\max_{u}\left[u\beta-\mathcal{E}(n,u)\right],$ (9)
where the relationship between $\beta$ and $u$ is given by the expected utility under the optimal control
$u=\mathcal{U}^{\dagger}(n,\beta):=\mathbb{E}_{k^{\dagger}(\cdot)}\left[\mathcal{U}(n(\cdot))\right]=\frac{\partial}{\partial\beta}V(n,\beta).$ (10)
See Appendix A for the derivation. Note that the formulation can be easily extended to the case with multiple constraints, e.g., $\mathbb{E}_{k(\cdot)}\left[\mathcal{U}_{j}(n(\cdot))|n(0)=n\right]\geq u_{j}$ for $j=1,\ldots,J$ or the equality constraint $\mathbb{E}_{k(\cdot)}\left[\mathcal{U}(n(\cdot))|n(0)=n\right]=u$. In the following, we focus on solving the unconstrained problem, Eq. (7).
We will calculate the optimal reaction rate coefficient $(k^{\dagger}(t))_{0\leq t<T}$ that attains the optimum $V(n_{0},\beta)$. The standard protocol for solving the problem is to consider the value function $V_{t}(n,\beta)$, which is defined as the maximum of the expectation in Eq. (7) from $t$ to $T$ under the condition $n(t)=n$:
@@
where $\mathcal{U}_{t}(n(\cdot))$ and $\mathcal{C}_{t}(n(\cdot),k(\cdot))$ are the utility and control cost functions from time $t$ to $T$, e.g., $\mathcal{U}_{t}(n(\cdot)):=U_{T}(n(T))+\int_{t}^{T}U_{\tau}(n(\tau))d\tau$. The value function is equal to the terminal utility function at time $T$, $V_{T}(n,\beta)=\beta U_{T}(n)$, and also provides the optimum for the original problem at time 0, $V_{0}(n,\beta)=V(n,\beta)$. Thus, the reaction rate attaining Eq. (11) is identical to the optimal reaction rates $k^{\dagger}$ in the original control problem.
To obtain the value function, we usually apply the dynamic programming principle to derive the Hamilton–Jacobi–Bellman (HJB) equation, a system of differential equations that $V_{t}(n,\beta)$ satisfies: for $t\in(0,T)$,
$\frac{\partial}{\partial t}V_{t}(n,\beta)=\beta U_{t}(n)+\max_{k\in\mathbb{R}_{\geq 0}^{|R|}}\left[\mathcal{L}_{k}V_{t}(n,\beta)-C(n,k)\right],$ (12)
with the terminal condition $V_{T}(n,\beta)=\beta U_{T}(n)$. $\mathcal{L}_{k}$ is the Kolmogorov backward operator (generator), which is given by, for any scalar function $f(n)$,
$\mathcal{L}_{k}f(n)=\sum_{r\in R}k_{r}h_{r}(n)\overline{\nabla}_{s_{r}}f(n),$ (13)
where $\overline{\nabla}_{s_{r}}f(n):=f(n+s_{r})-f(n)$ is the discrete directional derivative of $f$ along $s_{r}$. The optimal reaction rate coefficient is given by
$k^{\dagger}(t,n,\beta)=\arg\max_{k\in\mathbb{R}_{\geq 0}^{|R|}}\left[\mathcal{L}_{k}V_{t}(n,\beta)-C(n,k)\right],$ (14)
which explicitly depends on the current time $t$, state $n$, and $\beta$.
II.3 First exit optimal control problems
We have focused on the problem with a finite and fixed terminal time $T$. One can extend our approach to the problem with infinite time length $T\rightarrow\infty$. Since a stochastic reaction system often has absorbing states $\mathcal{N}_{\mathrm{abs}}\subset\mathcal{N}$, the stochastic process often reaches one of these absorbing states after some time $T_{\mathrm{exit}}:=\min\{t\geq 0|n(t)\in\mathcal{N}_{\mathrm{abs}}\}$. For example, in the birth-death processes
the extinction state $n_{A}=0$ is absorbing. The extinction events are particularly important in biological problems because the extinction of some species in a chemical, organismal, or human population could alter the dynamics of the system qualitatively. Thus, controlling species into either survival or extinction has many applications, as mentioned in Introduction. Moreover, biological control problems may not always have a prescribed end time $T$ because the goal of control is usually to achieve something rather than to do something until $T$. Thus, the first exit control is more essential than fixed-time control for biological problems.
We can consider the optimal control problem with such a random terminal time $T_{\mathrm{exit}}$. The utility function is replaced with the sum of the terminal utility function $U_{\mathrm{exit}}(n)$ defined on the absorbing states $\mathcal{N}_{\mathrm{abs}}$, and the instantaneous utility function $U(n)$ defined on the non-absorbing states $\mathcal{N}_{\mathrm{non}}:=\mathcal{N}\backslash\mathcal{N}_{\mathrm{abs}}$, i.e.,
$\mathcal{U}(n(\cdot))=U_{\mathrm{exit}}(n(T_{\mathrm{exit}}))+\int_{0}^{T_{\mathrm{exit}}}U(n(\tau))d\tau.$ (16)
Then, the value function becomes time-independent $V(n,\beta)=V_{t}(n,\beta)$, and the HJB equation can be derived as
$0=\beta U(n)+\max_{k\in\mathbb{R}_{\geq 0}^{|R|}}\left[\mathcal{L}_{k}V(n,\beta)-C(n,k)\right],$ (17)
with the boundary condition $V(n,\beta)=\beta U_{\mathrm{exit}}(n)$ for $n\in\mathcal{N}_{\mathrm{abs}}$. Once we have the value function $V(n,\beta)$, the optimal reaction rate is given by
$k^{\dagger}(n,\beta)=\arg\max_{k\in\mathbb{R}_{\geq 0}^{|R|}}\left[\mathcal{L}_{k}V(n,\beta)-C(n,k)\right],$ (18)
which depends on the current state $n$ and $\beta$ but not time $t$.
We will not delve into details in this work but we briefly discuss the formulation in Appendix D.
II.4 Choice of the control cost function
@@
To obtain the value function, we usually leverage the Hamilton–Jacobi–Bellman (HJB) equation, a differential equation that $V_{t}(n,\beta)$ satisfies. However, solving the HJB equation is generally intractable both analytically and even numerically because it is a nonlinear differential equation on a possibly infinite domain. This difficulty is the major obstacle to optimal control in applications. To address the difficulty, people are conventionally forced to restrict or approximate the original problem to a linear-quadratic-Gaussian (LQG) problem, in which only the linear dynamics on a continuous Euclidean space with Gaussian noise and a quadratic control cost function can be considered. However, all these restrictions of the LQG problem conflict with the essential properties of stochastic RN control, i.e., the nonlinearity of $h_{r}(n)$, the discreteness and nonnegativity of state $n\in\mathbb{Z}_{\geq 0}^{|X|}$, the nonnegativity of the control parameter $k_{r}\geq 0$, and Poissonian nature of stochasticity. Several studies attempted to overcome some of these difficulties [32, 33, 34, 35, 36, 37]. Nonetheless, optimal control was not practical for biological problems described by RN.
In this work, we demonstrate that the difficulty can be resolved by designing the control cost function appropriately for RN. The natural way to measure control costs would vary depending on the context. However, we find some basic assumptions that may be applied to the control cost function in most practical situations. Noting the similarity of these assumptions to the properties of the divergence, which is the dissimilarity measure between a pair of probability measures or nonnegative measures in information theory, we propose using the divergence to formulate the control cost.
First, the control cost $C$ should be a function of the increment in reaction counts $\xi_{r}$ or the reaction flux $j_{r}=\frac{d}{dt}\mathbb{E}[\xi_{r}]=k_{r}h_{r}(n)$, by which the control cost function depends on the state $n$ and control input $k$. This is a natural assumption, for instance, when the control of the reaction network is implemented by intermittent interventions activated at every chance of a reaction event.
Second, the instantaneous control cost function $C(n,k)$ should have a unique minimizer $k^{0}(n)=\arg\min_{k}C(n,k)$ for each $n\in\mathcal{N}$. Without loss of generality, the minimum control cost can be set to be zero, $C(n,k^{0}(n))=0$. Note that in conventional control problems such as LQG problems, the control cost function is given by a quadratic function, which is minimized when the control input is zero, $k^{0}_{r}=0$. However, the non-zero (positive) uncontrolled input $k^{0}_{r}>0$ is common in controlling reaction networks, for instance, when the controller can suppress reactions that otherwise lead to undesirable outcomes.
Figure 2: Diagram showing the properties of the control cost. The control of reaction fluxes is visualized as the control of flow speed in channels. $j_{r}^{0}$ and $j_{r}$ are the uncontrolled (input) and the controlled (output) reaction fluxes, respectively. $a,b\in[0,1]$ are the ratios of dividing the input and output fluxes.
The dynamics with vanishing control cost can be interpreted as the uncontrolled dynamics. Thus, the control cost function $C(n,k)$ measures the deviation of the current control input $k$ from the uncontrolled one $k^{0}(n)$. Combined with the first assumption, the control cost function can be interpreted as the control of reaction flux from the uncontrolled one $j_{r}^{0}(n):=k^{0}_{r}(n)h_{r}(n)$ to the controlled one $j_{r}$ (Fig. 2 (a)). Noting that the reaction fluxes $j$ and $j^{0}$ can be seen as nonnegative measures on the set of reactions $R$, the divergence in information theory may serve as a control cost function.
Specifically, we propose to use $f$-diverngence-like function
$C(n,k)=\sum_{r\in R}j^{0}_{r}f_{r}\left(\frac{j_{r}}{j^{0}_{r}}\right)=\sum_{r\in R}h_{r}(n)k^{0}_{r}f_{r}\left(\frac{k_{r}}{k^{0}_{r}}\right),$ (20)
where $f_{r}:[0,+\infty)\rightarrow[0,+\infty]$ is a convex function satisfying $f_{r}(1)=0$ and $f_{r}^{\prime}(1)=0$ for $r\in R$. Note that the standard definition of the $f$-divergence in the information theory [22, 24] adopts a reaction-independent function $f_{r}=f$ in Eq. (20) to ensure permutation invariance, but we extend it to allow reaction dependence. For example, using a quadratic function $f_{r}(x)=\frac{1}{2}(x-1)^{2}$, we obtain $\chi^{2}$-divergence-like control cost function
$C_{quad}(n,k)=\sum_{r\in R}h_{r}(n)\frac{\left(k_{r}-k^{0}_{r}\right)^{2}}{2k_{r}^{0}}.$ (21)
Quadratic cost functions weighted by the propensity $h_{r}(n)$ have been used in quantifying the biological cost of cellular reactions [38].
In fact, $f$-divergence is the only class of divergences with (1) decomposability, (2) subadditivity, and (3) positive homogeneity, all of which are reasonable assumptions as the control cost. The decomposability means that the total control cost is a sum of the control cost for each reaction $r\in R$ so that the control cost function can be written as $C=\sum_{r\in R}D_{r}(j_{r},j^{0}_{r})$, which implies the independence of the control mechanisms or flow-controlled channels for each reaction. The subadditivity and positive homogeneity mean that
$D_{r}(j_{r},j^{0}_{r})\leq D_{r}(aj_{r},bj^{0}_{r})+D_{r}((1-a)j_{r},(1-b)j^{0}_{r}),$ (22)
$D_{r}(\rho j_{r},\rho j^{0}_{r})=\rho D_{r}(j_{r},j^{0}_{r}),$ (23)
holds, respectively, for any $j_{r},j^{0}_{r},\rho\geq 0$ and $a,b\in[0,1]$. The subadditivity implies that the control cost increases if the input and output fluxes of a channel are divided into those of two independent channels (Fig. 2 (b)). However, positive homogeneity implies that if the division ratios for input and output fluxes are the same, $a=b$, the two channels are identical, and control can be done efficiently without increasing the control cost (Fig. 2 (c)). In Appendix B, we prove that the class of control cost functions satisfying the above assumptions is exactly $f$-divergence-like functions in Eq. (20).
Furthermore, provided an admissible control exists, we can guarantee the existence of a unique optimal reaction rate coefficient by imposing 1-coercivity and strict convexity on the function $f_{r}$. The 1-coercivity, i.e.,
$\lim_{x\rightarrow+\infty}\frac{f_{r}(x)}{x}=+\infty,$ (24)
and the strict convexity, i.e.,
$f_{r}(ax+(1-a)x^{\prime})<af_{r}(x)+(1-a)f_{r}(x^{\prime})$ (25)
for any $a\in(0,1)$ and $x,x^{\prime}\geq 0$, guarantee the existence and uniqueness of the optimum, respectively.
Using the $f$-divergence-like control cost function, the HJB Eq. (12) can be rewritten as
$\frac{\partial}{\partial t}V_{t}(n,\beta)=\beta U_{t}(n)+\sum_{r\in R}k^{0}_{r}h_{r}(n)f_{r}^{*}(\overline{\nabla}_{s_{r}}V_{t}(n,\beta)),$ (26)
where $f_{r}^{*}(y):=\max_{x}[xy-f_{r}(x)]$ is the Legendre transform of $f_{r}(x)$. The optimal reaction rate coefficient $k^{\dagger}_{r}(t,n,\beta)$ is given by the following time-dependent state-feedback controller
$k^{\dagger}_{r}(t,n,\beta)=k^{0}_{r}\partial f_{r}^{*}(\overline{\nabla}_{s_{r}}V_{t}(n,\beta)).$ (27)
As for the first exit problems, HJB Eq. (17) becomes
$0=\beta U(n)+\sum_{r\in R}k_{r}^{0}h_{r}(n)f_{r}^{*}(\overline{\nabla}_{s_{r}}V(n,\beta)),$ (28)
and the optimal reaction rate is given by
$k^{\dagger}_{r}(n,\beta)=k^{0}_{r}\partial f_{r}^{*}(\overline{\nabla}_{s_{r}}V(n,\beta)).$ (29)
In the case of quadratic control cost, the Legendre transform of $f_{r}(x)=\frac{1}{2}(x-1)^{2}$ is given by
$f^{*}_{r}(y)=\frac{1}{2}\left([y+1]_{+}^{2}-1\right),\quad\partial f^{*}_{r}(y)=[y+1]_{+},$ (30)
where $a]_{+}=\max\{a,0\}$ is the positive part of a scalar $a$. The positive part operation $\cdot$$_{+}$ comes from the non-negativity constraint on $k$.
Therefore, the $f$-divergence-like control cost functions are intuitive and mathematically convenient in formulating stochastic RN control problems.
II.5 Optimal control with KL cost
Among a variety of $f$-divergence-like control cost functions, we found that the following Kullback–Leibler (KL) divergence-like control cost function is exceptionally useful:
$C_{KL}(n,k)=\sum_{r\in R}h_{r}(n)\left(k_{r}\log\frac{k_{r}}{k_{r}^{0}}-k_{r}+k_{r}^{0}\right),$ (31)
which is given by setting $f_{r}(x)=x\log x-x+1$ in Eq. (20). The value at $x=0$ is defined by the limit $f_{r}(0):=\lim_{x\rightarrow 0}f_{r}(x)=1$.
The optimal reaction coefficient for $r\in R$ at time $t$ is given by